OpenStudy (anonymous):

Hello Please help with this problem. Lost on the steps. Simplifying Radicals: (m^1/4 n^1/2)^2 (m^2 n^3)^1/2

OpenStudy (jdoe0001):

do you know what to do when the base is the same and exponents differ?

OpenStudy (anonymous):

not really

OpenStudy (jdoe0001):

http://www.biology.arizona.edu/biomath/tutorials/exponential/graphics/Law1.gif <--- you'd need for this one, the rule 3, that is, distribute the outer exponent and rule 1

OpenStudy (anonymous):

okay I noticed the rules 1 and 3

OpenStudy (anonymous):

would I use rule first

OpenStudy (anonymous):

m*n=mn m*n=m^2 n^3

OpenStudy (jdoe0001):

http://mathb.in/12789

OpenStudy (anonymous):

Not sure how to use that site

OpenStudy (jdoe0001):

hmm ?

OpenStudy (jdoe0001):

do you see the right-hand side part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did you do all that with the codes

OpenStudy (jdoe0001):

hehehe, well, I use an editor... so :)

OpenStudy (jdoe0001):

so you see, same base, different exponents, rule 1 listed there then you distribute the expression's exponenet

OpenStudy (anonymous):

so would I end up with this m*n=mn m*n=m^2 n^3

OpenStudy (jdoe0001):

hmm is that what you've got?

OpenStudy (anonymous):

wait I forgot to square root

OpenStudy (anonymous):

square root m n square root m^2 n^3

OpenStudy (jdoe0001):

hmmm so did you see the expression at the link?

OpenStudy (anonymous):

which link

OpenStudy (jdoe0001):

http://mathb.in/12789 <---

OpenStudy (anonymous):

1/4m (9/4)n(7/2) following the link there I think this is the answer

OpenStudy (jdoe0001):

http://mathb.in/12792 <--- look closely

OpenStudy (jdoe0001):

the rule 1 from http://www.biology.arizona.edu/biomath/tutorials/exponential/graphics/Law1.gif is used firstly then rule 3

OpenStudy (anonymous):

Okay, thank you I will work on studying this more. You have been a lot of help

OpenStudy (jdoe0001):

yw